# leftmargin = 20pt, rightmargin = 20pt, backgroundcolor = bggray, middlelinecolor = black, roundcorner = 15

### v1

@@ Abstract
The NOSTR is a communication protocol for the social web, based on the w3c websockets standard. Although it is still in its infancy, it is well known as a social media protocol, thousands of trusted users and multiple user interfaces, offering a unique experience and enormous capabilities. To name a few, the NOSTR applications include but are not limited to direct messaging, file sharing, audio/video streaming, collaborative writing, blogging and data processing through distributed AI directories. In this work, we propose an approach that builds upon the existing protocol structure with end goal a decentralized marketplace for federated learning and LLM training. In this proposed design there are two parties: on one side there are customers who provide a dataset that they want to use for training an AI model. On the other side, there are service providers, who receive (parts of) the dataset, train the AI model, and for a payment as an exchange, they return the optimized AI model. The decentralized and censorship resistant features of the NOSTR enable the possibility of designing a fair and open marketplace for training AI models and LLMs.
@@ 1 Introduction
In parallel with the NOSTR protocol development, distributed AI training has attracted major attention the last few years, with the Language Model breakthroughs leading the progress. By combining recent advanced tools and breakthroughs from open communication protocols [1], bitcoin payments [7] and distributed optimization algorithms [8], we propose a system design that enables open marketplaces for training AI models. Specifically, the NOSTR has already built-in protocol flow for payments through the lightning network [7], as well as communication flows for a distributed data processing system, which makes it the state of the art protocol for designing a decentralized market for distributed computation and AI directories. Such a network system can mitigate privacy issues, security risks and potentially minimize operational costs of machine learning and language models training by leveraging the structure and capabilities of the NOSTR protocol. We begin by discussing some fundamental aspects of the NOSTR, and the core protocol application designed specifically for the social web.
A prominent challenge for social web applications is that of designing decentralized social networks through peer-to-peer architectures. Additionally, web 3.0 and blockchain implementations will either fail to scale or they will be highly centralized. Tokenized web 3.0 approaches introduce unnecessarily highly counter-party economic risk. Although decentralized payment systems require both blockchain and Proof-of-Work [9, 10] (or smart contracts) for asynchronous distributed consensus under faulty processes, it becomes more and more clear that other applications including decentralized social media and marketplaces for computational resources work securely and efficiently without blockchain or global consensus. In contrast, with peer-to-peer networks, web 3.0 and blockchain applications, the NOSTR protocol works efficiently as a social media protocol, because it does not require a majority of nodes of the network to share a common database of the data history. As a consequence, it is also economically efficient since it does not require any form of tokens and it does not hide any counter party risk. In summary, the NOSTR offers a censorship-resistance, open, public network structure, as a well as a digital IDs (DIDs) solution through cryptography and a novel protocol implementation [11].
@@
Users do not communicate directly with each other, but rather through relays and relays interact with users only. Relays listen only and propagate information. Although relays can not change the content (or action) of users, they can possibly act independently (with any other user or relay) and refuse to propagate pieces of information. The protocol is robust to such censorship attempts, since anyone including users can run one or multiple relays and users can choose an arbitrary set of relays to interact with. Also anyone can run their own relay. Clients retrieve and publish data from relays of their choice. As an example, assume that the User A "follows" a set S of Users, then the client just queries selected relays for posts from the public key of all users within the set S. That is, on startup a client queries data from all relays it knows for all users it follows (for example, all updates from the last day), then displays that data to the user chronologically. An event (user action, for instance a post or a like) can contain any kind of structured data, and clients choose how to display events (defined as any user generated action) of certain kind, while relays can handle them seamlessly. Examples of different kinds are simple note/post (kind ), re-post (kind ), reaction (kind ) [1]. The NOSTR protocol offers flexible structure of messages through a customizable structure and different kind of events.
@@ 2 Related Work
The orchestration of middleware systems to distribute computational tasks over wide-ranging networks marks a cornerstone in the evolution of distributed computing. A pivotal advancement in this area has been the development of grid computing. Emerging as a potent global cyber-infrastructure, grid computing is engineered to bolster the forthcoming wave of e-Science applications. By amalgamating large-scale, diverse, and distributed resources, it fosters a cooperative model for the sharing of computational power and data storage on an unprecedented scale. This cooperative paradigm paves the way for unparalleled computational and data management capabilities, enabling the execution of sophisticated scientific, engineering, and academic endeavors previously beyond reach.
@@
Within the diverse ecosystem of grid computing frameworks, centralized systems such as Pegasus [14] and Triana[15] have emerged as paragons. Pegasus, functioning as a workflow mapping engine, abstracts workflow execution from its underlying infrastructure, facilitating the seamless amalgamation of computational tasks across varied grid environments. This system is heralded for its ability to streamline and optimize complex workflows with remarkable efficiency. Conversely, Triana offers a dynamic platform for the composition of applications, harnessing distributed resources through meticulous integration and orchestration. These frameworks underscore the dynamic and adaptable nature of grid computing, illustrating its capacity to tackle complex challenges through the strategic leverage of distributed resources.
@@
Concurrently, volunteer computing has established itself as an integral component within the broader spectrum of grid systems. The Berkeley Open Infrastructure for Network Computing (BOINC)[16] stands as a testament to this model, capturing the essence of volunteer computing by mobilizing the latent processing power of personal computers worldwide for scientific research necessitating extensive computational resources. This model champions the democratization of scientific inquiry, inviting public engagement while significantly enhancing computational prowess without the need for equivalent infrastructural investment.
@@
The "BOINC: A Platform for Volunteer Computing" [17] discourse elucidates the enduring success of the BOINC project, spotlighting its sustained operation over the years as a beacon of collaborative scientific endeavor. This analysis brings to the forefront the prohibitive costs and reliability concerns associated with conventional cloud computing platforms, such as Amazon Web Services (AWS), which, despite their widespread adoption, pose financial and operational challenges for extensive, computation-intensive projects. Furthermore, the discourse addresses the intrinsic challenge of maintaining a robust volunteer base — a critical component for the sustenance of volunteer computing frameworks. The fluctuating levels of engagement and the logistical complexities of coordinating such a decentralized workforce underscore the necessity for innovative solutions to bolster volunteer retention and participation.
@@
To address the lack of incentives and participation in volunteer computing, reward based approaches have been introduced in prior works. For instance, Gridcoin[18] is a proof-of-stake cryptocurrency that rewards participants for contributing computational power to scientific research projects, notably through the Berkeley Open Infrastructure for Network Computing (BOINC) and Folding@Home. However, a proof-of-stake cryptocurrency may introduce several technical issues, since the payment network requires a global synchronized clock [10]. Such issues and significant outages become prominent in many non-proof-of-work blockchain systems (for instance [19]).
@@
While blockchain based approaches aim to provide innovative solutions to incentivize contributions in distributed computing, such approaches differ significantly from the NOSTR protocol. The NOSTR offers a broader set of capabilities, extending beyond the computational resource sharing model to include secure communications and permissionsless interaction between customers and service providers. In contrast with volunteer computing and proof-of-stake cryptocurency models, the decentralized protocol NOSTR integrates bitcoin through the lightning network and a dedicated protocol flow for payments. The NOSTR solves the user-to-user micro-payment problem, incentivizes honest collaborative computation, enables the possibility of an open market through instant, permissionless payments and presents a viable solution to the challenge of volunteer maintenance. These together with the secure and robust digital IDs implementation, make the NOSTR the state of the art protocol, capable of supporting a wide range of decentralized applications and services, offering a pathway to enhance viability, sustainability, and effectiveness.
@@ 3 A Decentralized Marketplace for Distributed Computing
As the name of the protocol indicates (Notes and Other Stuff Transmitted by Relays), the type of communication messages and user-to-user interactions are not restricted to simple notes, and they support a variety of alternative applications including direct messaging, file sharing, audio/video streaming, collaborative writing, marketplaces for data processing and more [20]. Another example is that of marketplaces for data processing, where users interact with each other to execute AI algorithms. For the implementation of data processing on the NOSTR protocol, dedicated event kinds have been developed to facilitate distributed data processing (see [21, 22]). These implementations have been considered for designing a decentralized AI-directories marketplace of the form "Money-In Data-Out". For instance, a customer can provide some input data, such as text or audio, and pay a service provider to receive a short summary of the input data. More specifically, the kinds in range [21] are reserved for text manipulation jobs including text extraction, summarising text, text generation and translation. Image generation corresponds to kind , and video/audio manipulation corresponds to kinds in range .
In this work, we propose an extension of the protocol. Our goal is to design a decentralized marketplace for federated learning and LLMs training. In a nutshell, a customer provides a data-set, model specification and a required payment to a set of service providers. Then the service providers execute the model training, and they return the parameters of the model in order to receive the payment for their work (Money-In AI-Out). In other words, we propose a protocol flow for developing AI-Model Vending Machines (AI-VM), where a customer provides a data-set, a set of specifications, and submits a payment to a set of service providers to acquire a model trained on the input data-set. There are several components of the protocol that allow us to build upon existing protocol rules and design a decentralized marketplace. These include Digital IDs, payments, fast communication though websockets, a flexible protocol structure, also suitable for distributing computing. We proceed by briefly explaining these features.
@@ 4 Signatures & Digital IDs
The NOSTR identity is a dual-key cryptographic system [13, 23, Schnorr Signatures]. Each user profile is associated with a private (nsec) and a public key (npub). These keys are generated locally without the need of a trusted party. Cryptographic keys are essential to the protocol security, ensuring the protection of user identities and messages. The private key is used to sign messages and other actions by the user, verifying that they originate from the authentic identity owner. On the other hand, the public key is employed to authenticate these messages, confirming that they were signed using the corresponding private key. Identically, to the social media applications, the same DID, private and public keys can be used for alternative applications, for instance private messaging applications, or marketplaces. This allows for a consistent and secure way to authenticate users DIDs and across different applications over the NOSTR protocol.
@@
Owing a well known DID or building a solid reputation over the NOSTR offers the benefits of recognizability and trust from other users, similarly to all other online media platforms. Notably, this identity and its accompanying reputation can be seamlessly transferred between different applications. In the context of a marketplace, a good reputation can help a service provider maximize their profit. A high reputation indicates a consistent service, which in turn makes more customers trust the service provider, ensuring a steady stream of business. Alternatively, bad reputation may assist customers to avoid incompetent service providers. As a matter of fact, a reputation system for a decentralized and open marketplace for distributing computing will be valuable. However, this remains out of the scope of this work, but it is an interesting implementation possibility for future extensions of the protocol.
@@ 5 Protocol Design for Distributed Computing and AI Directories
The NOSTR protocol supports a primal version of AI directories for data processing in the form of an open and decentralized marketplace [21, 22], known as Data Vending Machines. On one side, there are users (customers) who provide some data that should be processed (image, video or text) and they request jobs that can be done with some AI tools, for instance image generation from text, a transcript of a video, a summary of an article. On the other side, there are AI agents (service providers) that offer their services to users, and in exchange for a payment, they deliver the requested results in the form of processed data. This system works in the form of Money-In Data-out; a marketplace for data processing/manipulation by applying AI tools. This design sets the foundations of distributed data processing over the NOSTR protocol and offers a unique protocol structure and flow for additional development.
Our goal is to utilize the existing architecture and to go one step further by extending components of the protocol if necessary, in order to design a decentralized marketplace for Federated Learning (FL) and Large Language Models (LLM) training. In what we propose later (see Section 7), users (customers) provide an input data-set and certain specifications and request a sequence of jobs that aim to train a deep neural network or an LLM model. Then service providers receive the data, model specifications and the requested task (train through regular FL or LLM training (e.g., via the DiLoCo algorithm) [8]), they train the model through multiple rounds and jobs requests, and they return the model parameters. In exchange for their computational resources, the service providers receive recurrent payments from the customer upon delivering valid results. A main difference of our design in comparison with existing approaches is the concept of delivering a trained model instead of processed data; a set of service providers are required to optimize the same model in parallel and deliver valid results through multiple rounds of computation. For that purpose we extend the structure of protocol by introducing new event kinds, a new protocol flow and we deploy the classical FL approach, as well as the DiLoCo algorithm [8] for LLM training. Next, we delve into the existing protocol structure for distributing computing over the NOSTR network. We also emphasize the modifications we introduce to support the Money-IN AI-Out marketplace design.
@@ 5.1 Distributed Computing on the NOSTR Protocol
Herein, we discuss the existing dedicated protocol messages and protocol flow for distributed computing. We mainly focus on five components that we will utilize later with slight extensions or variations, these include job request events, job result events, job feedback events, events for service provider discoverability and job chaining rules. These type of events consist the core of the communication messages between customer and service providers. We refer the reader to the NOSTR Implementation Possibility - 90 (NIP-90 [24]) for further details.
Basic Protocol Flow for Data Vending Machines (DMVs) [24]
@@
Service Providers may submit job-feedback events (kind:7000, e.g. payment-required, processing, error, etc.).
@@
At any point, if there is an amount pending to be paid as instructed by the service provider, the user can pay the included bolt11 invoice [25] of the job result event that the service provider has sent to the user.
The above protocol is designed to be deliberately ambiguous. For instance a customer may choose a set of service providers and restrict or not restrict the job request to be completed by this specific set of service providers. Also, a service provider may not start a job until they receive a form of payment, or their response can depend on the other participants reputation based on their public key. The flexibility of the protocol flow allows us to extend parts of it, and implement additional communication steps for the design of on demand LLMs training over a decentralized marketplace. Next we proceed by presenting the existing events as appear in the protocol flow for DMVs.
@@
A Job Request for Data Vending Mchines is an event, published by a customer through a set of relays. This event signals that a customer is interested in receiving the result of some kind of computation by an AI tool. The reserved range for job request events is . The field with label "content" is empty, while, all the required information should be provided in the field with label "tags". The field "tags" contains all the required information from an AI-agent in order to process the data. {mdframed}
For completeness, we provide additional explanation regarding each tag of the events. We first present the existing events structure. Then we introduce any necessary changes or additions that we will consider in the protocol flow and in the algorithm. For the Job Request event (Event Type 1) there are the following tags: {outline}\1 <i>: All the required input data for the job appear here. Specifically, the inputs of the AI algorithm. \2 <data>: The argument for the input. \2 <input-type>: The type of the argument input, it must be one of the following: \3 url: A URL to be fetched of the data that should be processed. \3 event: A Nostr event ID. \3 job: The output of a previous job with the specified event ID. The determination of which output to build upon is up to the service provider to decide (e.g. waiting for a signaling from the customer, waiting for a payment, etc.) \3 text: <data> is the value of the input, no resolution is needed \1 <relay>: The relay where the event/job was published \1 <marker>: An optional field indicating how this input should be used within the context of the job \1 <output>: Expected output format. \1 <param>: Optional parameters for the job. Different job request kind defines this more precisely. (e.g. [ "task", "run option", "initial state" ]). \1 <bid>: Customer MAY specify a maximum amount (in millisats) they are willing to pay \1 <relays>: List of relays where Service Providers SHOULD publish responses to. \1 p: Service providers the customer is interested in. Other service providers might still choose to process the job.
@@
The existing job request events include a large number of different jobs, dedicated to data processing applications. Such job request have a certain structure (see Event Type 1). For clarity and for efficacy of our approach, we propose to introduce the kinds 8000 - 8999 as Job Request events for federated learning, LLM training and distributed optimization. Although, each kind can correspond at a different algorithm or variation of an algorithm, specific details can be considered by introducing additional fields in the tags. As an example, notice that the predefined structure of Event Type 1, requires a single input data type as we discussed earlier, while additional information may be included in parameters field (<param>).
Similarly, for the purpose of training an LLM, the customer needs to provide a training data-set though a URL as input data, and additionally a set of parameters related to the training task, for instance initial model parameters, a URL to the source code of the training or optimization method, model specification (example "LLaMA-2"), rules for validation of the output. To maintain the existing protocol flow as much as possible, we consider the input data as a URL to the training dataset for the initial job request, or the output of a previous job with a specified event job id to declare a job chaining and multiple rounds of training. In both cases (initial or chain job request), we consider all the required information for executing algorithm in the parameters fields for consistency. Below some pieces of that information follow (Event Type 2):
a source for the code: While the service provider may have its own code, there should be reference to a standard approach that guarantees consistency
@@
job details: expected execution type or hardware that is needed, rules related to time-out conditions.
@@
Through the optimization process we introduce a new job request at each round. We also consider job chaining to speed up the process and most likely to execute the training by assigning the jobs to the same service providers, as long as they provide valid outputs. A running job will be assigned to new service provider, if a previous service provider goes offline, or fails to provide valid result. Finally, the customer has the option to encrypt all the parameters, and only the selected service provider(s) in the tag "p" will be able to decrypt them. We proceed by briefly discussing the Job Chaining rules, for a detailed description we refer the reader to [24, (NIP-90)].
@@
A Customer has the option to request multiple rounds of jobs to be processed as a chain, where the output of a job is the input of another job. In the context of data processing and DMVs this can be a sequence of processing as a podcast transcription by one service provider and then extracting a summary of the transcription by another. One way to implement this is by specifying as input an event id of a different job with the job type. In general, service providers may begin processing a subsequent job the moment they see the prior job result, but they will likely wait to receive a (partial) payment first.
In the case of federated learning and LLM training such options still exist, however we propose certain protocol rules to minimize the risk introduced by bad actors and malicious users. To do this, we consider necessary for the customer to validate the progress made by each service provider at every single round. After the validation is completed, each service provider who produced successful results receives a partial payment for the work done in the corresponding round. Then the customer updates the current state of model and submits a follow-up job request for the next round (and optionally a partial payment to the service providers upfront). Upon completion of the optimization task at each round, the service providers communicate with the customer through a Job Result event.

### v2

# A Decentralized Marketplace for Federated Learning and LLM Training on the NOSTR Protocol

@@ Abstract
The NOSTR is a communication protocol for the social web, based on the w3c websockets standard. Although it is still in its infancy, it is well known as a social media protocol, with thousands of trusted users and multiple user interfaces, offering a unique experience and enormous capabilities. To name a few, the NOSTR applications include but are not limited to direct messaging, file sharing, audio/video streaming, collaborative writing, blogging and data processing through distributed AI directories. In this work, we propose an approach that builds upon the existing protocol structure with end goal a decentralized marketplace for federated learning and LLM training. In this proposed design there are two parties: on one side there are customers who provide a dataset that they want to use for training an AI model. On the other side, there are service providers, who receive (parts of) the dataset, train the AI model, and for a payment as an exchange, they return the optimized AI model. To demonstrate viability, we present a proof-of-concept implementation over public NOSTR relays. The decentralized and censorship resistant features of the NOSTR enable the possibility of designing a fair and open marketplace for training AI models and LLMs.
@@ 1 Introduction
In parallel with the NOSTR protocol development, distributed AI training has attracted major attention over the last few years, with the Language Model breakthroughs leading the progress. By combining recent advanced tools and breakthroughs from open communication protocols [1], bitcoin payments [7] and distributed optimization algorithms [8], we propose a system design that enables open marketplaces for training AI models. Specifically, the NOSTR has already built-in protocol flow for payments through the lightning network [7], as well as communication flows for a distributed data processing system, which makes it the state of the art protocol for designing a decentralized market for distributed computation and AI directories. Such a network system can mitigate privacy issues, security risks and potentially minimize operational costs of machine learning and language models training by leveraging the structure and capabilities of the NOSTR protocol. We begin by discussing some fundamental aspects of the NOSTR, and the core protocol application designed specifically for the social web.
A prominent challenge for social web applications is designing decentralized social networks through peer-to-peer architectures. Additionally, web 3.0 and blockchain implementations will either fail to scale or they will be highly centralized. Tokenized web 3.0 approaches introduce unnecessarily highly counter-party economic risk. While decentralized payment systems require both blockchain and Proof-of-Work [9, 10] (or smart contracts) for asynchronous distributed consensus under faulty processes, it becomes increasingly clear that other applications including decentralized social media and marketplaces for computational resources work securely and efficiently without blockchain or global consensus. In contrast to peer-to-peer networks, web 3.0 and blockchain applications, the NOSTR protocol works efficiently as a social media protocol, because it does not require a majority of nodes of the network to share a common database of the data history. As a consequence, it is also economically efficient since it does not require any form of tokens and it does not hide any counter party risk. In summary, the NOSTR offers a censorship-resistance, open, public network structure, as a well as a digital IDs (DIDs) solution through cryptography and a novel protocol implementation [11].
@@
Users do not communicate directly with each other, but rather through relays and relays interact with users only. Relays passively listen and propagate information. Although relays can not change the content (or action) of users, they can act independently with any other user or relay and refuse to propagate pieces of information. The protocol is robust to such censorship attempts, since anyone including users can run one or multiple relays and users can choose an arbitrary set of relays to interact with. Also anyone can run their own relay. Clients retrieve and publish data from relays of their choice. As an example, assume that the User A "follows" a set S of Users, then the client just queries selected relays for posts from the public key of all users within the set S. That is, on startup a client queries data from all relays it knows for all users it follows (for example, all updates from the last day), then displays that data to the user chronologically. An event (user action, for instance a post or a like) can contain any kind of structured data, and clients choose how to display events (defined as any user generated action) of certain kind, while relays can handle them seamlessly. Examples of different kinds are simple note/post (kind ), re-post (kind ), reaction (kind ) [1]. The NOSTR protocol offers flexible structure of messages through a customizable structure and different kind of events.
@@ 2 Related Work
The orchestration of middleware systems to distribute computational tasks over wide-ranging networks is a cornerstone in the evolution of distributed computing. A pivotal advancement in this area has been the development of grid computing. Emerging as a potent global cyber-infrastructure, grid computing is engineered to bolster the forthcoming wave of e-Science applications. By integrating large-scale, diverse, and distributed resources, grid computing fosters a cooperative model for the sharing of computational power and data storage on an unprecedented scale. This cooperative paradigm paves the way for unparalleled computational and data management capabilities, enabling the execution of sophisticated scientific, engineering, and academic endeavors previously beyond reach.
@@
Within the grid computing ecosystem, centralized systems such as Pegasus [14] and Triana[15] have emerged as paragons. Pegasus, functioning as a workflow mapping engine, abstracts workflow execution from its underlying infrastructure, facilitating the seamless amalgamation of computational tasks across varied grid environments. This system is heralded for its ability to streamline and optimize complex workflows with remarkable efficiency. Conversely, Triana offers a dynamic platform for the composition of applications, harnessing distributed resources through meticulous integration and orchestration. These frameworks underscore the dynamic and adaptable nature of grid computing, illustrating its capacity to tackle complex challenges through the strategic leverage of distributed resources.
@@
Concurrently, volunteer computing has established itself as an integral component within the broader spectrum of grid systems. The Berkeley Open Infrastructure for Network Computing (BOINC)[16] exemplifies this model, capturing the essence of volunteer computing by mobilizing the latent processing power of personal computers worldwide for scientific research necessitating extensive computational resources. This model promotes the democratization of scientific inquiry, inviting public engagement while significantly enhancing computational prowess without the need for equivalent infrastructural investment.
@@
The "BOINC: A Platform for Volunteer Computing" [17] highlights the enduring success of the BOINC project, demonstrating its sustained operation over the years as a beacon of collaborative scientific endeavor. This analysis brings to the forefront the prohibitive costs and reliability concerns associated with conventional cloud computing platforms, such as Amazon Web Services (AWS), which, despite their widespread adoption, pose financial and operational challenges for extensive, computation-intensive projects. Furthermore, the discourse addresses the intrinsic challenge of maintaining a robust volunteer base — a critical component for the sustenance of volunteer computing frameworks. The fluctuating levels of engagement and the logistical complexities of coordinating such a decentralized workforce underscore the necessity for innovative solutions to bolster volunteer retention and participation.
@@
To address the lack of incentives and participation in volunteer computing, reward-based approaches have been introduced in prior works. For instance, Gridcoin[18] is a proof-of-stake cryptocurrency that rewards participants for contributing computational power to scientific research projects, notably through the Berkeley Open Infrastructure for Network Computing (BOINC) and Folding@Home. However, a proof-of-stake cryptocurrency may introduce several technical issues, because the payment network requires a global synchronized clock [10]. Such issues and significant outages become prominent in many non-proof-of-work blockchain systems (for instance [19]).
@@
While blockchain based approaches aim to provide innovative solutions to incentivize contributions in distributed computing, such approaches differ significantly from the NOSTR protocol. NOSTR offers a broader set of capabilities, extending beyond the computational resource sharing model to include secure communications and permissionsless interaction between customers and service providers. In contrast with volunteer computing and proof-of-stake cryptocurrency models, the decentralized protocol NOSTR integrates bitcoin through the lightning network and a dedicated protocol flow for payments. NOSTR solves the user-to-user micro-payment problem, incentivizes honest collaborative computation, enables an open market through instant, permissionless payments and presents a viable solution to the challenge of volunteer maintenance. These, together with the secure and robust digital IDs implementation, make NOSTR the state-of-the-art protocol, capable of supporting a wide range of decentralized applications and services, offering a pathway to enhance viability, sustainability, and effectiveness.
@@ 3 A Decentralized Marketplace for Distributed Computing
As the name of the protocol indicates (Notes and Other Stuff Transmitted by Relays), the type of communication messages and user-to-user interactions are not restricted to simple notes, and they support a variety of alternative applications including direct messaging, file sharing, audio/video streaming, collaborative writing, marketplaces for data processing and more [20]. Another example is marketplaces for data processing, where users interact with each other to execute AI algorithms. For the implementation of data processing on the NOSTR protocol, dedicated event kinds have been developed to facilitate distributed data processing (see [21, 22]). These implementations have enabled the design of a decentralized AI directory marketplace of the form "Money-In Data-Out". For instance, a customer can provide some input data, such as text or audio, and pay a service provider to receive a short summary of the input data. More specifically, the kinds in range [21] are reserved for text manipulation jobs including text extraction, summarising text, text generation and translation. Image generation corresponds to kind , and video/audio manipulation corresponds to kinds in range .
In this work, we propose an extension of the protocol. Our goal is to design a decentralized marketplace for federated learning and LLMs training. In a nutshell, a customer provides a dataset, model specification and a required payment to a set of service providers. Then the service providers execute the model training, and they return the parameters of the model in order to receive the payment for their work (Money-In AI-Out). In other words, we propose a protocol flow for developing AI-Model Vending Machines (AI-VM), where a customer provides a dataset, a set of specifications, and submits a payment to a set of service providers to acquire a model trained on the input dataset. There are several components of the protocol that allow us to build upon existing protocol rules and design a decentralized marketplace. These include Digital IDs, payments, fast communication though websockets, a flexible protocol structure, also suitable for distributing computing. We proceed by briefly explaining these features.
@@ 4 Signatures & Digital IDs
The NOSTR identity is a dual-key cryptographic system [13, 23, Schnorr Signatures]. Each user profile is associated with a private (nsec) and a public key (npub). These keys are generated locally without the need of a trusted party. Cryptographic keys are essential to the protocol’s security, ensuring the protection of user identities and messages. The private key is used to sign messages and other actions by the user, verifying that they originate from the authentic identity owner. On the other hand, the public key is employed to authenticate these messages, confirming that they were signed using the corresponding private key. Similarly to social media applications, the same DID, private and public keys can be used for alternative applications, for instance private messaging applications, or marketplaces. This allows for a consistent and secure way to authenticate users DIDs across different applications over the NOSTR protocol.
@@
Owning a well-known DID or building a solid reputation over the NOSTR offers the benefits of recognizability and trust from other users, similarly to all other online media platforms. Notably, this identity and its accompanying reputation can be seamlessly transferred between different applications. In the context of a marketplace, a good reputation can help service providers maximize their profit. A high reputation indicates a consistent service, which in turn makes more customers trust the service provider, ensuring a steady stream of business. Alternatively, bad reputation may assist customers to avoid incompetent service providers. As a matter of fact, a reputation system for a decentralized and open marketplace for distributing computing will be valuable. However, this remains outside the scope of this work, but it is an interesting implementation possibility for future extensions of the protocol.
@@ 5 Protocol Design for Distributed Computing and AI Directories
The NOSTR protocol supports a preliminary version of AI directories for data processing in the form of an open and decentralized marketplace [21, 22], known as Data Vending Machines. On one side, there are users (customers) who provide some data that should be processed (image, video or text) and they request jobs that can be done with some AI tools, for instance image generation from text, a transcript of a video, a summary of an article. On the other side, there are AI agents (service providers) that offer their services to users, and in exchange for a payment, they deliver the requested results in the form of processed data. This system works in the form of Money-In Data-Out; a marketplace for data processing/manipulation by applying AI tools. This design sets the foundations of distributed data processing over the NOSTR protocol and offers a unique protocol structure and flow for additional development.
Our goal is to utilize the existing architecture and to go one step further by extending components of the protocol if necessary, in order to design a decentralized marketplace for Federated Learning (FL) and Large Language Models (LLM) training. In what we propose later (see Section 7), users (customers) provide an input dataset and certain specifications and request a sequence of jobs that aim to train a deep neural network or an LLM model. Then service providers receive the data, model specifications and the requested task (train through regular FL or LLM training (e.g., via the DiLoCo algorithm) [8]), they train the model through multiple rounds and jobs requests, and they return the model parameters. In exchange for their computational resources, the service providers receive recurrent payments from the customer upon delivering valid results. The main difference between our design and existing approaches is the concept of delivering a trained model instead of processed data; a set of service providers are required to optimize the same model in parallel and deliver valid results through multiple rounds of computation. For that purpose we extend the structure of the protocol by introducing new event kinds, a new protocol flow and we deploy the classical FL approach, as well as the DiLoCo algorithm [8] for LLM training. Next, we delve into the existing protocol structure for distributing computing over the NOSTR network. We also emphasize the modifications we introduce to support the Money-IN AI-Out marketplace design.
@@ 5.1 Distributed Computing on the NOSTR Protocol
Herein, we discuss the existing dedicated protocol messages and protocol flow for distributed computing. We mainly focus on five components that we will utilize later with slight extensions or variations, these include job request events, job result events, job feedback events, events for service provider discoverability, and job chaining rules. These type of events consist the core of the communication messages between customer and service providers. We refer the reader to the NOSTR Implementation Possibility - 90 (NIP-90 [24]) for further details.
Basic Protocol Flow for Data Vending Machines (DVMs) [24]
@@
Service Providers may submit job feedback events (kind:7000, e.g. payment-required, processing, error, etc.).
@@
At any point, if a payment is pending as instructed by the service provider, the user can pay the bolt11 invoice [25] included in the corresponding job result event.
The protocol is designed to be intentionally flexible. For instance, a customer may choose a set of service providers and restrict or not restrict the job request to be completed by this specific set of service providers. Also, a service provider may not start a job until they receive a form of payment, or their response can depend on the other participants reputation based on their public key. The flexibility of the protocol flow allows us to extend parts of it, and implement additional communication steps for the design of on demand LLMs training over a decentralized marketplace. Next we proceed by presenting the existing events as appear in the protocol flow for DVMs.
@@
A Job Request for Data Vending Mchines is an event, published by a customer through a set of relays. This event signals that a customer is interested in receiving the result of some kind of computation by an AI tool. The reserved range for job request events is . The field with label "content" is empty, while, all the required information should be provided in the field with label "tags". The field "tags" contains all the required information from an AI-agent in order to process the data.
"kind": 5xxx // kind in range 5000 - 5999,
"content": "",
"tags":
Event Type 1: Job Request Event with kind in range 5000 -5999
For completeness, we provide additional explanation regarding each tag of the events. We first present the existing events structure. Then we introduce any necessary changes or additions that we will consider in the protocol flow and in the algorithm. For the Job Request event (Event Type 1) there are the following tags:
<i>: All the required input data for the job appear here. Specifically, the inputs of the AI algorithm.
<data>: The argument for the input.
<input-type>: The type of the argument input, it must be one of the following:
url: A URL to be fetched of the data that should be processed.
event: A Nostr event ID.
job: The output of a previous job with the specified event ID. The determination of which output to build upon is up to the service provider to decide (e.g. waiting for a signaling from the customer, waiting for a payment, etc.)
text: <data> is the value of the input, no resolution is needed
<relay>: The relay where the event/job was published
<marker>: An optional field indicating how this input should be used within the context of the job
<output>: Expected output format.
<param>: Optional parameters for the job. Different job request kind defines this more precisely. (e.g. [ "task", "run option", "initial state" ]).
<bid>: Customer MAY specify a maximum amount (in millisats) they are willing to pay
<relays>: List of relays where Service Providers SHOULD publish responses to.
p: Service providers the customer is interested in. Other service providers might still choose to process the job.
@@
The existing job request events support a large number of different jobs, dedicated to data processing applications. Such job request have a certain structure (see Event Type 1). For clarity and the efficacy of our approach, we propose to introduce the kinds 8000 - 8999 as Job Request events for federated learning, LLM training and distributed optimization. Although each kind can correspond to a different algorithm or variation of an algorithm, specific details can be considered by introducing additional fields in the tags. As an example, notice that the predefined structure of Event Type 1, requires a single input data type as we discussed earlier, while additional information may be included in parameters field (<param>).
"kind": 8xxx // kind in range 8000 - 8999,
"content": "",
"tags":
Event Type 2: Job Request Event for training AI models, kind in range 8000 -8999
Similarly, for the purpose of training an LLM, the customer needs to provide a training dataset through a URL as input data, and additionally a set of parameters related to the training task, for instance, initial model parameters, a URL to the source code of the training or optimization method, model specification (example "LLaMA-2"), rules for validation of the output. To maintain the existing protocol flow as much as possible, we consider the input data as a URL to the training dataset for the initial job request, or the output of a previous job with a specified job event ID to indicate a job chaining and multiple rounds of training. In both cases (initial or chain job request), we consider all the required information for algorithm execution in the parameters fields for consistency. Below some pieces of that information follow (Event Type 2):
a source for the code: While the service provider may have its own code, there should be a reference to a standard approach that guarantees consistency
@@
job details: expected execution time or required hardware, rules related to timeout conditions.
@@
Through the optimization process, we introduce a new job request at each round. We also consider job chaining to speed up the process and most likely to execute the training by assigning the jobs to the same service providers, as long as they provide valid outputs. A running job will be reassigned to a new service provider if a previous service provider goes offline, or fails to provide valid result. Finally, the customer can encrypt all parameters so that only the selected service provider(s) in tag "p" can decrypt them. We proceed by briefly discussing the Job Chaining rules, for a detailed description we refer the reader to [24, (NIP-90)].
@@
A customer has the option to request multiple rounds of jobs to be processed as a chain, where the output of a job is the input of another job. In the context of data processing and DVMs this can be a sequence of processing as a podcast transcription by one service provider and then extracting a summary of the transcription by another. One way to implement this is by specifying an event ID of a previous job as input, using the job type field. In general, service providers may begin processing a subsequent job the moment they see the prior job result, but they will likely wait to receive a (partial) payment first.
In the case of federated learning and LLM training such options still exist; however, we propose certain protocol rules to minimize the risk introduced by bad actors and malicious users. To do this, we consider it necessary for the customer to validate the progress made by each service provider at every single round. After the validation is completed, each service provider who produced successful results receives a partial payment for the work done in the corresponding round. Then the customer updates the current state of the model and submits a follow-up job request for the next round (and optionally a partial payment to the service providers upfront). Upon completion of the optimization task at each round, the service providers communicate with the customer through a Job Result event.